OpenStudy (anonymous):

A manufacturer produces two models of toy airplanes. It takes the manufacturer 20 minutes to assemble model A and 10 minutes to package it. It takes the manufacturer 15 minutes to assemble model B and 12 minutes to package it. In a given week, the total available time for assembling is 1800 minutes, and the total available time for packaging is 1080 minutes. Model A earns a profit of $10 for each unit sold and model B earns a profit of $8 for each unit sold. Assuming the manufacturer is able to sell as many models as it makes, how many units of each model should be produced to maximize the pro

OpenStudy (anonymous):

*profit for the given week?

OpenStudy (anonymous):

PLEASE HELP

OpenStudy (anonymous):

I got 90 for Model A and 0 for Model B

OpenStudy (anonymous):

What section are you in... systems of linear inequalities?

OpenStudy (anonymous):

Yes, I think so. I'm in an online class and usually it shows the topic up top but it doesn't for this word problem.

OpenStudy (anonymous):

Let x = the number of model A airplanes Let y = the number of model B airplanes Total profit = 10x + 8y ---

OpenStudy (anonymous):

Yeah I got 90 of Model A and 0 of Model B using that formula but that just doesn't seem right at all.

OpenStudy (anonymous):

Assembly time = 30x + 15y Packaging time = 10x + 12y Total time = 40x + 27y or 1080 = 40x + 27y

OpenStudy (anonymous):

oops... 20 minutes... Assembly time = 20x + 15y Packaging time = 10x + 12y Total time = 30x + 27y or 1080 = 30x + 27y

OpenStudy (anonymous):

Wait would 30 and 27 be the answers to the word problem?

OpenStudy (anonymous):

No 30 those are times in minutes... give me a few minutes.

OpenStudy (anonymous):

Okay thank you

OpenStudy (anonymous):

" total available time for assembling is 1800 minutes, and the total available time for packaging is 1080 minutes." ok... 1800 = 20x + 15y <--- assembly time 1080 = 10x + 12y <--- packaging time Solve this system... http://www.wolframalpha.com/input/?i=solve {1800+%3D+20x+%2B+15y%2C+1080+%3D+10x+%2B+12y} x = 60 model A's and y = 40 model B's Not sure about this because I did not use the profit information??

OpenStudy (anonymous):

Alright thanks for trying it though!

OpenStudy (anonymous):

Linear programming! Check this link... is this what you are doing? http://www.purplemath.com/modules/linprog.htm 1800 < 20x + 15y 1080 < 10x + 12y 0 < 10x + 8y <--- profit greater than zero ---